# Post-Minkowskian Theory Meets the Spinning Effective-One-Body Approach for Bound-Orbit Waveforms

### v1

@@ Abstract
Introduction – Since the initial detection of a gravitational wave (GW) from a binary–black-hole (BBH) merger Abbott et al. (2016a), the LIGO-Virgo-KAGRA (LVK) Collaboration Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021) and independent analyses have identified about 100 mergers of compact binaries Abbott et al. (2017a, 2023a); Nitz et al. (2023); Olsen et al. (2022); Mehta et al. (2023); Wadekar et al. (2023). These observations have begun to reveal the distributions of BH-masses and spins Abbott et al. (2023b), improved constraints on the neutron-star (NS) equation of state Abbott et al. (2018a), obtained independent measurements of the Hubble-Lemaître parameter Abbott et al. (2017b, 2023c), and validated General Relativity (GR) Abbott et al. (2016b, 2021a, 2021b).
@@
Enhancements in the sensitivity of current GW detectors, coupled with the development of next-generation (XG) observatories like the Einstein Telescope and Cosmic Explorer Punturo et al. (2010); Kalogera et al. (2021); Reitze et al. (2019), as well as future space-based detectors such as LISA Amaro-Seoane et al. (2017), TianQin Luo et al. (2016) or Taiji Wu et al. (2021), are poised to dramatically increase the number of detectable GW sources. These advancements will enable observations with a signal-to-noise ratio up to two orders of magnitude higher than what is currently achievable Borhanian and Sathyaprakash (2022), necessitating a commensurate improvement in the accuracy of waveform models. Recent research Dhani et al. (2024) has demonstrated that even state-of-the-art waveform models, designed for quasi-circular, spin-precessing BBHs, exhibit systematic biases when applied to future LVK runs and XG detectors. This bias becomes pronounced, especially for high-spin rates and significant asymmetries in spins and masses. Addressing the challenge of waveform accuracy is essential to realizing the full scientific potential of future runs and detectors Abbott et al. (2018b); Evans et al. (2021); Bogdanov et al. (2022); Borhanian and Sathyaprakash (2022); Colpi et al. (2024), and avoiding false claims of GR violations Toubiana et al. (2024); Gupta et al. (2024).
@@
Waveform models for compact binaries are crafted by synergistically combining analytical and numerical relativity (NR) results. NR tackles the formidable task of solving Einstein’s equations on supercomputers Pretorius (2005); Campanelli et al. (2006); Baker et al. (2006), a process notorious for its time-intensive nature. On the other hand, perturbative methods are used to obtain approximate solutions to the Einstein’s equations, offering analytic formulas that are swift to evaluate. Three primary perturbative approaches have been developed: (i) post-Newtonian (PN) theory Futamase and Itoh (2007); Blanchet (2014); Porto (2016); Schäfer, Gerhard and Jaranowski, Piotr (2018); Levi (2020); Jaranowski and Schaefer (1998); Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2014); Jaranowski and Schäfer (2015); Bernard et al. (2016, 2017); Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2016); Blanchet et al. (2023a, b); Foffa and Sturani (2019); Foffa et al. (2019); Blümlein, J. and Maier, A. and Marquard, P. and Schäfer, G. (2020) applicable in the weak-field and small-velocity limit, post-Minkowskian (PM) theory Westpfahl and Goller (1979); Westpfahl and Hoyler (1980); Bel et al. (1981); Westpfahl (1985); Schäfer (1986); Ledvinka et al. (2008); Buonanno et al. (2022); Travaglini et al. (2022); Bjerrum-Bohr et al. (2022); Kosower et al. (2022) in the weak-field regime, and the gravitational-self force (GSF) formalism Mino et al. (1997); Quinn and Wald (1997); Barack et al. (2002); Barack and Ori (2003); Gralla and Wald (2008); Detweiler (2008); Keidl et al. (2010); van de Meent (2018); Pound (2012); Pound et al. (2020); Gralla and Lobo (2022); Pound and Wardell (2021); Warburton et al. (2021); Wardell et al. (2023) for the small mass-ratio limit. These analytical results are then synthesized in the effective-one-body (EOB) approach Buonanno and Damour (1999, 2000); Damour et al. (2000); Damour (2001); Buonanno et al. (2006), which efficiently resums the perturbative calculations for the inspiral while retaining known non-perturbative results for BHs, achieving high accuracy for current observing runs via calibration to NR Buonanno et al. (2007a, b); Damour and Nagar (2008); Pan et al. (2011a); Damour et al. (2013); Pan et al. (2014); Taracchini et al. (2014); Bohé et al. (2017); Babak et al. (2017); Nagar et al. (2018); Ossokine et al. (2020); Gamba et al. (2022); Nagar et al. (2023); Pompili et al. (2023); Ramos-Buades et al. (2023).
@@
Thus far, the EOB waveform models utilized by the LVK Collaboration  [^1] The LVK Collaboration also utilizes phenomenological frequency- and time-domain waveform models, which are built by combining PN, EOB, and NR results Ajith et al. (2007); Pratten et al. (2021); Estellés et al. (2022). Additionally, it employs waveforms directly interpolated from NR simulations Blackman et al. (2017b); Varma et al. (2019b), where available. have primarily relied on resummations of the PN expansion, with the exception of Refs. Ramos-Buades et al. (2023); van de Meent et al. (2023), which included second-order GSF results Warburton et al. (2021) for the gravitational modes and radiation-reaction force. Given recent advancements in PM Damour (2016); Cheung et al. (2018); Bern et al. (2019a, 2023); Kosower et al. (2019); Bjerrum-Bohr et al. (2018); Cristofoli et al. (2019); Damgaard et al. (2019); Brandhuber et al. (2021); Vines (2018); Kälin, Gregor and Porto, Rafael A. (2020); Kälin, Gregor and Liu, Zhengwen and Porto, Rafael A. (2020); Mogull et al. (2021); Jakobsen et al. (2022a, b) and GSF Warburton et al. (2021); Wardell et al. (2023), there is now significant interest in exploring and developing waveform models that combine information from various perturbative methods in innovative ways. The aim is to address the waveform-accuracy challenge. In this regard, the PM approach is particularly interesting since an (n + 1) PM-order Hamiltonian includes all information up to the nPN order, and additional weak-field/high-velocity information from infinitely higher PN orders, making it suitable for systems with high velocities or large eccentricities at fixed periastron distances Khalil et al. (2022). Using sophisticated quantum-field-theory–based methods, tremendous progress has been made on the precision PM-frontier using both scattering amplitudes Bern et al. (2021a, 2022a); Damgaard et al. (2023) and worldlines Dlapa et al. (2022, 2023a); Jakobsen et al. (2023a, 2024). This progress is largely due to a blend of a clever and efficient organization of perturbative calculations and formal mathematical developments in understanding the properties of multi-loop integrals Parra-Martinez et al. (2020); Weinzierl (2022); Abreu et al. (2022); Blümlein and Schneider (2022); Frellesvig et al. (2024a, b); Klemm et al. (2024). These developments were primarily driven over the last several decades to address precision-collider physics. Furthermore, several sophisticated techniques, such as generalized unitarity Bern et al. (1995); Bern and Huang (2011), double copy Bern et al. (2010, 2019b, 2022b); Adamo et al. (2022), supersymmetry Jakobsen et al. (2022a) and massive higher spins Chung et al. (2019); Arkani-Hamed et al. (2020); Bern et al. (2021b); Bautista et al. (2023a, b); Chiodaroli et al. (2022); Aoude et al. (2023); Cangemi et al. (2023) have also been used to further enhance these computations
@@
In this paper, leveraging on Refs. Khalil et al. (2023); Pompili et al. (2023); Buonanno et al. (2024), we present the first PM-informed spinning EOB waveform model: SEOBNR-PM, encompassing the inspiral, as well as, the merger and ringdown phases. This model incorporates the most recent findings from PM theory into the EOB Hamiltonian, and it is mildly calibrated to NR waveforms. The SEOB-PM Hamiltonian (so-named as it does not include an NR calibration term) includes the non-spinning (conservative) 4PM Bern et al. (2022a); Dlapa et al. (2022) and spinning 5PM terms Guevara et al. (2019); Bern et al. (2021b); Kosmopoulos and Luna (2021); Chen et al. (2022); Jakobsen and Mogull (2022, 2023); Febres Cordero et al. (2023); Jakobsen et al. (2023a), alongside the known non-spinning 4PN Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2014); Jaranowski and Schäfer (2015); Bernard et al. (2016, 2017); Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2016) contributions, which also corrects the tails from unbound to bound orbits up to that order  [^2] We plan to update our SEOBNR-PM model in the near future to include the recent findings in Ref. Dlapa et al. (2024), which allow to derive the local-in-time EOB Hamiltonian for bound orbits at 4PM.. Our PM counting is a physical one, with spin orders contributing in addition to loop orders (see Table II in Ref. Buonanno et al. (2024)). We construct our SEOBNR-PM model within the pySEOBNR code Mihaylov et al. (2023), which was recently built to make the development of SEOBNR models, including the calibration to NR waveforms, highly efficient. As an example, we show in Fig. 1 the agreement between the (mildly) calibrated SEOBNR-PM and NR for a spinning BBH coalescence.
@@
The EOB framework for waveforms – We use geometric units $G=1=c$, and set $M=m_{1}+m_{2}$ and $\nu=m_{1}m_{2}/M^{2}$, where $m_{1}$ and $m_{2}$ are the BH’s masses. In the EOB formalism, the binary’s conservative dynamics is described by the EOB Hamiltonian $H_{\rm EOB}=M\,\sqrt{1+2\nu(H_{\rm eff}/\mu-1)}$, where $H_{\rm eff}$ is the Hamiltonian of an effective test-body of mass $\mu=\nu\,M$ moving in the (deformed) Kerr spacetime, being $40\leq\nu\leq 1/$ 4 the deformation parameter. We also introduce the mass ratio $q=m_{1}/m_{2}>1$. We limit to nonprecessing spins (i.e., aligned spins) and introduce the spin lengths $a_{i}=m_{i}\chi_{i}$, with $a_{\pm}=M\chi_{\pm}=a_{1}\pm a_{2}$.
In the center-of-masss frame, the inspiral-plunge dynamics, for aligned-spin BHs, is computed from the EOB equations of motion Buonanno and Damour (2000); Pan et al. (2011a):
@@
where ($r$,$\phi$,$p_{r}$,$p_{\phi}$) are the polar-coordinates canonical variables. (The construction of $H_{\rm eff}$ will be described in the next section.) Employing results from the SEOBNRv5 model Khalil et al. (2023); Pompili et al. (2023), the radiation-reaction force ($\mathcal{F}_{\phi}$) is computed by summing over the PN GW modes (augmented with GSF information van de Meent et al. (2023)) in a factorized form Damour et al. (2009); Damour and Nagar (2007); Pan et al. (2011b); Damour and Nagar (2008); Pompili et al. (2023), which are used to obtain the inspiral-plunge modes after enhancing them by non-quasi-circular corrections Damour et al. (2003); Taracchini et al. (2014); Bohé et al. (2017); Pompili et al. (2023) during the plunge.
@@
Following Ref. Buonanno et al. (2024), the $\nu$-corrections with respect to the probe limit are built into the $A$-potential and the gyro-gravitomagnetic factors as $A={(1-2u+\chi_{+}^{2}u^{2}+\Delta A)}/{[1+\chi_{+}^{2}u^{2}(2u+1)]}$ and $g_{a_{\pm}}={\Delta g_{a_{\pm}}/}{u^{2}}$. These respectively carry the even- and odd-in-spin corrections, and are PM-expanded up to a physical 5PM ($u^{5}$) order (see Table II in Ref. Buonanno et al. (2024)):
@@
An important subtlety within our Hamiltonian is the presence of non-local-in-time contributions (tails). These imply a dependence on the full past history of the binary, and thus distinguish between elliptic and hyperbolic (scattering) trajectories. In the scattering Hamiltonian presented in Ref. Buonanno et al. (2024), tails are signaled by factors of $\log(\gamma^{2}-1)$, which develops an imaginary part when $\gamma<1$ for bound orbits. To produce a real Hamiltonian, we therefore replace $\log(\gamma^{2}-1)\to\log(u)$ (see Supplemental Material for details). We also include the 4PN non-spinning bound-orbit correction $\Delta A^{\rm 4PN}$ in Eq. (8):
@@
ensuring the correct bound-orbit dynamics at 4PN order in the non-spinning case (the numerical coefficients $c_{i}$ are provided in the Supplemental Material). We verify our complete EOB Hamiltonian up to 4.5PN order Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2014); Antonelli et al. (2020); Mandal et al. (2023); Kim et al. (2023); Levi and Steinhoff (2015) by finding a suitable canonical transformation to its PN-expanded counterpart  [^3] We thank Mohammed Khalil for providing us with suitable PN Hamiltonians to compare with.. The non-spinning component is determined only up to quadratic order in eccentricity ($p_{r}^{2}$) in the tail integral, as higher powers in eccentricity appear at lower-PM orders. Thus, we ensured that the 1PM–3PM (tail-free) non-spinning dynamics are unmodified by the presence of the 4PN correction (5444 As a spin-dependent bound-orbit correction would only arise at 5.5PN order, we choose not to include it. .
@@
Figure 3: Spin-orbit (left panel) and spin-squared (right panel) contributions to the binding energy, for an equal-mass BBH, as a function of the (quasi-circular) velocity $v=(M\dot{\phi})^{2/3}$, for the (calibrated) SEOBNRv5 with ($a_{6}$,$d_{\rm SO}$) and (uncalibrated) SEOB-PM Hamiltonians at different PM orders. The vertical line represents the merger of the NR configuration (the one at the lowest velocity among those used), with the number of GW cycles (top axis) referring to the same simulation. The shaded regions are estimates of the NR uncertainty Ossokine et al. (2018). The lower panel shows the absolute value of the fractional difference. The feature in the lower-right panel around $v\sim 0.4$ is due to a zero-crossing.
Finally, let us comment on the appearance of special functions in $H_{\rm eff}$. Starting at 3PM order we encounter the combination ${{\rm arccosh}(\gamma)}/{\sqrt{\gamma^{2}-1}}$. As ${\rm arccosh}(\gamma)$ and $\sqrt{1-\gamma^{2}}$ are both imaginary-valued when $\gamma<1$, we find it convenient to replace this combination by ${\arccos(\gamma)}/{\sqrt{1-\gamma^{2}}}$, which has the same small-velocity expansion for scattering kinematics. At 4PM order we then encounter logarithms, dilogarithms (${\rm Li}_{2}$) and elliptic functions (K/E) of the first and second kind, all evaluated as functions of $\gamma$. In this case, we also find it convenient to introduce the inverse tangent integral ${\rm Ti}_{2}(x):=\int_{0}^{x}({\rm d}t/t)\arctan t$, analogously to what is done above. Fast numerical routines exist for evaluating all of these functions in Cython Behnel et al. (2011), and this leads to an efficient numerical evaluation within pySEOBNR Mihaylov et al. (2023).
@@
We also extract different spin contributions to the binding energy by combining results from NR simulations for various equal-mass spin combinations Dietrich et al. (2017); Ossokine et al. (2018): ${\cal{E}}_{\mathrm{SO}}=-\frac{1}{6}(-0.6,0)+\frac{8}{3}(0.3,0)-2(0,0)-\frac{1}{2}(0.6,0)+\mathcal{O}(S^{3})$ and ${\cal{E}}_{\mathrm{S^{2}}}=\frac{3}{2}(-0.6,0)-2(0,0)+\frac{3}{2}(0.6,0)-(0.6,-0.6)+\mathcal{O}(S^{3})$, where the numbers in brackets correspond to the dimensionless spins $(\chi_{1},\chi_{2})$ of the BHs. In Fig. 3 we illustrate the spin-orbit and spin-squared contributions for an equal-mass BBH to the binding energy for the (uncalibrated) SEOB-PM at different PM orders, as compared with NR and with the (calibrated) SEOBNRv5 with ($a_{6}$, $d_{\rm SO}$). Despite not being calibrated to NR, SEOB-PM shows excellent agreement with the NR results, with a clear convergence towards the NR prediction, as more PM corrections are included. Its accuracy is somewhat better than SEOBNRv5, despite the latter model using a Hamiltonian calibrated in the non-spinning ($a_{6}$) and spin-orbit coupling ($d_{\rm SO}$) sector (see below).

### v2

@@ Abstract
Introduction – Since the initial detection of a gravitational wave (GW) from a binary-black-hole (BBH) merger Abbott et al. (2016a), the LIGO-Virgo-KAGRA Collaboration Aasi et al. (2015); Acernese et al. (2015); Akutsu et al. (2021) and independent analyses have identified about 100 mergers of compact binaries Abbott et al. (2017a, 2023a); Nitz et al. (2023); Olsen et al. (2022); Mehta et al. (2023); Wadekar et al. (2023). These observations have begun to reveal the distributions of BH masses and spins Abbott et al. (2023b), improved constraints on the neutron-star equation of state Abbott et al. (2018a), obtained independent measurements of the Hubble-Lemaître parameter Abbott et al. (2017b, 2023c), and validated General Relativity Abbott et al. (2016b, 2021a, 2021b).
@@
Enhancements in the sensitivity of current GW detectors, coupled with the development of next-generation observatories like the Einstein Telescope and Cosmic Explorer Punturo et al. (2010); Kalogera et al. (2021); Reitze et al. (2019), as well as future space-based detectors such as LISA Amaro-Seoane et al. (2017), TianQin Luo et al. (2016) or Taiji Wu et al. (2021), are poised to dramatically increase the number of detectable GW sources. These advancements will enable observations with a signal-to-noise ratio up to two orders of magnitude higher than what is currently achievable Borhanian and Sathyaprakash (2024), necessitating a commensurate improvement in the accuracy of waveform models. Recent research Dhani et al. (2024) has demonstrated that even state-of-the-art waveform models, designed for quasi-circular, spin-precessing BBHs, exhibit systematic biases when applied to future LIGO-Virgo-KAGRA runs and next-generation detectors. This bias becomes pronounced, especially for high-spin rates and significant asymmetries in spins and masses. Addressing the challenge of waveform accuracy is essential to realizing the full scientific potential of future runs and detectors Abbott et al. (2018b); Evans et al. (2021); Bogdanov et al. (2022); Borhanian and Sathyaprakash (2024); Colpi et al. (2024) and avoiding false claims of General Relativity violations Toubiana et al. (2024); Gupta et al. (2024).
@@
Waveform models for compact binaries are crafted by synergistically combining analytical and numerical relativity (NR) results. NR tackles the formidable task of solving Einstein’s equations on supercomputers Pretorius (2005); Campanelli et al. (2006); Baker et al. (2006), a process notorious for its time-intensive nature. On the other hand, perturbative methods are used to obtain approximate solutions to Einstein’s equations, offering analytic formulas that are swift to evaluate. Three primary perturbative approaches have been developed: post-Newtonian (PN) theory Futamase and Itoh (2007); Blanchet (2014); Porto (2016); Schäfer, Gerhard and Jaranowski, Piotr (2018); Levi (2020); Jaranowski and Schaefer (1998); Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2014); Jaranowski and Schäfer (2015); Bernard et al. (2016, 2017); Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2016); Blanchet et al. (2023a, b); Foffa and Sturani (2019); Foffa et al. (2019); Blümlein, J. and Maier, A. and Marquard, P. and Schäfer, G. (2020) applicable in the weak-field and small-velocity limit, post-Minkowskian (PM) theory Westpfahl and Goller (1979); Westpfahl and Hoyler (1980); Bel et al. (1981); Westpfahl (1985); Schäfer (1986); Ledvinka et al. (2008); Buonanno et al. (2022); Travaglini et al. (2022); Bjerrum-Bohr et al. (2022); Kosower et al. (2022) in the weak-field regime, and the gravitational-self force (GSF) formalism Mino et al. (1997); Quinn and Wald (1997); Barack et al. (2002); Barack and Ori (2003); Gralla and Wald (2008); Detweiler (2008); Keidl et al. (2010); van de Meent (2018); Pound (2012); Pound et al. (2020); Gralla and Lobo (2022); Pound and Wardell (2021); Warburton et al. (2021); Wardell et al. (2023) for the small mass-ratio limit. These analytical results are then synthesized in the effective-one-body (EOB) approach Buonanno and Damour (1999, 2000); Damour et al. (2000); Damour (2001); Buonanno et al. (2006), which efficiently resums the perturbative calculations for the inspiral while retaining known non-perturbative results for BHs, achieving high accuracy for current observing runs via calibration to NR Buonanno et al. (2007a, b); Damour and Nagar (2008); Pan et al. (2011a); Damour et al. (2013); Pan et al. (2014); Taracchini et al. (2014); Bohé et al. (2017); Babak et al. (2017); Nagar et al. (2018); Ossokine et al. (2020); Gamba et al. (2022); Nagar et al. (2023); Pompili et al. (2023); Ramos-Buades et al. (2023).
@@
Thus far, the EOB waveform models utilized by the LIGO-Virgo-KAGRA Collaboration [^1] The LIGO-Virgo-KAGRA Collaboration also utilizes phenomenological frequency- and time-domain waveform models, which are built by combining PN, EOB, and NR results Ajith et al. (2007); Pratten et al. (2021); Estellés et al. (2022). Additionally, it employs waveforms directly interpolated from NR simulations Blackman et al. (2017a); Varma et al. (2019a), where available. have primarily relied on resummations of the PN expansion, with the exception of Refs. Ramos-Buades et al. (2023); van de Meent et al. (2023), which included second-order GSF results Warburton et al. (2021) for the gravitational modes and radiation-reaction force. Given recent advancements in PM Damour (2016); Cheung et al. (2018); Bern et al. (2019, 2023, 2024a); Kosower et al. (2019); Bjerrum-Bohr et al. (2018); Cristofoli et al. (2019); Damgaard et al. (2019); Brandhuber et al. (2021); Vines (2018); Kälin, Gregor and Porto, Rafael A. (2020); Kälin, Gregor and Liu, Zhengwen and Porto, Rafael A. (2020); Mogull et al. (2021); Jakobsen et al. (2022a, b); Bhattacharyya et al. (2024) and GSF Warburton et al. (2021); Wardell et al. (2023), there is now significant interest in exploring and developing waveform models that combine information from various perturbative methods in innovative ways. The aim is to address the waveform-accuracy challenge. In this regard, the PM approach is particularly interesting since an (n + 1) PM-order ($G^{n+1}$) Hamiltonian includes all information up to the nPN order (wherein $v^{2}/c^{2}\sim GM/(rc^{2})$), and additional weak-field, high-velocity information from infinitely higher PN orders, making it suitable for systems with high velocities or large eccentricities at fixed periastron distances Khalil et al. (2022). Using sophisticated quantum-field-theory-based methods, tremendous progress has been made on the precision PM frontier using both scattering amplitudes Bern et al. (2021a, 2022a); Damgaard et al. (2023) and worldlines Dlapa et al. (2022, 2023a); Jakobsen et al. (2023a, 2024). This progress is largely due to a blend of a clever and efficient organization of perturbative calculations and formal mathematical developments in understanding the properties of multi-loop integrals Parra-Martinez et al. (2020); Weinzierl (2022); Abreu et al. (2022); Blümlein and Schneider (2022); Frellesvig et al. (2024a, b); Klemm et al. (2024). These developments were primarily driven over the past several decades to address precision-collider physics. Furthermore, several sophisticated techniques, such as generalized unitarity Bern et al. (1995); Bern and Huang (2011), double copy Bern et al. (2010, 2024b, 2022b); Adamo et al. (2022), supersymmetry Jakobsen et al. (2022a) and massive higher spins Chung et al. (2019); Arkani-Hamed et al. (2020); Bern et al. (2021b); Bautista et al. (2023a, b); Chiodaroli et al. (2022); Aoude et al. (2023); Cangemi et al. (2024) have also been used to further enhance these computations.
@@
In this Letter, leveraging on Refs. Khalil et al. (2023); Pompili et al. (2023); Buonanno et al. (2024), we present the first PM-informed spinning EOB waveform model: SEOBNR-PM, encompassing the inspiral, as well as the merger and ringdown phases. This model incorporates the most recent findings from PM theory into the EOB Hamiltonian, and it is mildly calibrated to NR waveforms. The SEOB-PM Hamiltonian (so named as it does not include an NR calibration term) includes the non-spinning (conservative) 4PM Bern et al. (2022a); Dlapa et al. (2022) and spinning 5PM terms Guevara et al. (2019); Bern et al. (2021b); Kosmopoulos and Luna (2021); Chen et al. (2022); Jakobsen and Mogull (2022, 2023); Febres Cordero et al. (2023); Jakobsen et al. (2023a), alongside the known non-spinning 4PN Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2014); Jaranowski and Schäfer (2015); Bernard et al. (2016, 2017); Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2016) contributions, which also corrects the tails from unbound to bound orbits up to that order. [^2] We plan to update our SEOBNR-PM model in the near future to include the recent findings in Ref. Dlapa et al. (2024), which enables derivation of the local-in-time 4PM EOB Hamiltonian for bound orbits. Our PM counting is a physical one, with spin orders contributing in addition to loop orders (see Table II in Ref. Buonanno et al. (2024)). We construct our SEOBNR-PM model within the pySEOBNR code Mihaylov et al. (2023), which was recently built to make the development of SEOBNR models, including the calibration to NR waveforms, highly efficient. As an example, we show in Fig. 1 the agreement between the (mildly) calibrated SEOBNR-PM and NR for a spinning BBH coalescence.
@@
EOB framework for waveforms – We use geometric units $G=1=c$, and set $M=m_{1}+m_{2}$ and $\nu=m_{1}m_{2}/M^{2}$, where $m_{1}$ and $m_{2}$ are the BH’s masses. In the EOB formalism, the binary’s conservative dynamics is described by the EOB Hamiltonian $H_{\rm EOB}=M\,\sqrt{1+2\nu(H_{\rm eff}/\mu-1)}$, where $H_{\rm eff}$ is the Hamiltonian of an effective test body of mass $\mu=\nu\,M$ moving in the (deformed) Kerr spacetime, with $40\leq\nu\leq 1/$ 4 the deformation parameter. We also introduce the mass ratio $q=m_{1}/m_{2}>1$. We limit to nonprecessing spins (i.e., aligned spins) and introduce the spin lengths $a_{i}=m_{i}\chi_{i}$, with $a_{\pm}=M\chi_{\pm}=a_{1}\pm a_{2}$.
In the center-of-mass frame, the inspiral-plunge dynamics, for aligned-spin BHs, is computed from the EOB equations of motion Buonanno and Damour (2000); Pan et al. (2011a):
@@
where ($r$,$\phi$,$p_{r}$,$p_{\phi}$) are the canonical variables in polar coordinates. (The construction of $H_{\rm eff}$ will be described in the next section.) Employing results from the SEOBNRv5 model Khalil et al. (2023); Pompili et al. (2023), the radiation-reaction force ($\mathcal{F}_{\phi}$) is computed by summing over the PN GW modes (augmented with GSF information van de Meent et al. (2023)) in a factorized form Damour et al. (2009); Damour and Nagar (2007); Pan et al. (2011b); Damour and Nagar (2008); Pompili et al. (2023), which are used to obtain the inspiral-plunge modes after enhancing them by non-quasi-circular corrections Damour et al. (2003); Taracchini et al. (2014); Bohé et al. (2017); Pompili et al. (2023) during the plunge.
@@
Following Ref. Buonanno et al. (2024), the $\nu$-corrections with respect to the probe limit are built into the $A$-potential and the gyro-gravitomagnetic factors as $A={(1-2u+\chi_{+}^{2}u^{2}+\Delta A)}/{[1+\chi_{+}^{2}u^{2}(2u+1)]}$ and $g_{a_{\pm}}={\Delta g_{a_{\pm}}/}{u^{2}}$. These respectively carry the even- and odd-in-spin corrections, and are PM expanded up to a physical 5PM ($u^{5}$) order (see Table II in Ref. Buonanno et al. (2024)):
@@
An important subtlety within our Hamiltonian is the presence of non-local-in-time contributions (tails). These imply a dependence on the full past history of the binary, and thus distinguish between elliptic and hyperbolic (scattering) trajectories. In the scattering Hamiltonian presented in Ref. Buonanno et al. (2024), tails are signaled by factors of $\log(\gamma^{2}-1)$, which develops an imaginary part when $\gamma<1$ for bound orbits. To produce a real Hamiltonian, we therefore replace $\log(\gamma^{2}-1)\to\log(u)$ (see Supplemental Material for details). We also include the 4PN non-spinning bound-orbit correction $\Delta A^{\rm 4PN}$ in Eq. (4),
@@
ensuring the correct bound-orbit dynamics at 4PN order in the non-spinning case (the numerical coefficients $c_{i}$ are provided in the Supplemental Material). We verify our complete EOB Hamiltonian up to 4.5PN order Damour, Thibault and Jaranowski, Piotr and Schäfer, Gerhard (2014); Antonelli et al. (2020); Mandal et al. (2023); Kim et al. (2023); Levi and Steinhoff (2015) by finding a suitable canonical transformation to its PN-expanded counterpart. [^3] We thank Mohammed Khalil for providing us with suitable PN Hamiltonians to compare with. The non-spinning component is determined only up to quadratic order in eccentricity ($p_{r}^{2}$) in the tail integral, as higher powers in eccentricity appear at lower PM orders. Thus, we ensured that the 1PM–3PM (tail-free) non-spinning dynamics are unmodified by the presence of the 4PN correction (5). [^4] As a spin-dependent bound-orbit correction would only arise at 5.5PN order, we choose not to include it.
@@
Figure 3: Spin-orbit (left-hand panel) and spin-squared (right-hand panel) contributions to the binding energy, for an equal-mass BBH, as a function of the (quasi-circular) velocity $v=(M\dot{\phi})^{2/3}$, for the (calibrated) SEOBNRv5 with ($a_{6}$,$d_{\rm SO}$) and (uncalibrated) SEOB-PM Hamiltonians at different PM orders. The vertical line represents the merger of the NR configuration (the one at the lowest velocity among those used), with the number of GW cycles (top axis) referring to the same simulation. The shaded regions are estimates of the NR uncertainty Ossokine et al. (2018). The lower panel shows the absolute value of the fractional difference. The feature in the lower-right panel around $v\sim 0.4$ is due to a zero-crossing.
Finally, let us comment on the appearance of special functions in $H_{\rm eff}$. Starting at 3PM order we encounter the combination ${{\rm arccosh}(\gamma)}/{\sqrt{\gamma^{2}-1}}$. As ${\rm arccosh}(\gamma)$ and $\sqrt{1-\gamma^{2}}$ are both imaginary valued when $\gamma<1$, we find it convenient to replace this combination by ${\arccos(\gamma)}/{\sqrt{1-\gamma^{2}}}$, which has the same small-velocity expansion for scattering kinematics. At 4PM order we then encounter logarithms, dilogarithms (${\rm Li}_{2}$) and elliptic functions (K/E) of the first and second kind, all evaluated as functions of $\gamma$. In this case, we also find it convenient to introduce the inverse tangent integral ${\rm Ti}_{2}(x):=\int_{0}^{x}({\rm d}t/t)\arctan t$, analogously to what is done above. Fast numerical routines exist for evaluating all of these functions in Cython Behnel et al. (2011), and this leads to an efficient numerical evaluation within pySEOBNR Mihaylov et al. (2023).
@@
We also extract different spin contributions to the binding energy by combining results from NR simulations for various equal-mass spin combinations Dietrich et al. (2017); Ossokine et al. (2018): ${\cal{E}}_{\mathrm{SO}}=-\frac{1}{6}{\cal{E}}(-0.6,0)+\frac{8}{3}{\cal{E}}(0.3,0)-2{\cal{E}}(0,0)-\frac{1}{2}{\cal{E}}(0.6,0)+\mathcal{O}(S^{3})$ and ${\cal{E}}_{\mathrm{S^{2}}}=\frac{3}{2}{\cal{E}}(-0.6,0)-2{\cal{E}}(0,0)+\frac{3}{2}{\cal{E}}(0.6,0)-{\cal{E}}(0.6,-0.6)+\mathcal{O}(S^{3})$, where ${\cal{E}}(\chi_{1},\chi_{2})$ denotes the binding energy in a simulation with dimensionless spins $\chi_{i}$. In Fig. 3 we illustrate the spin-orbit and spin-squared contributions for an equal-mass BBH to the binding energy for the (uncalibrated) SEOB-PM at different PM orders, as compared with NR and with the (calibrated) SEOBNRv5 with ($a_{6}$, $d_{\rm SO}$). Despite not being calibrated to NR, SEOB-PM shows excellent agreement with the NR results, with a clear convergence toward the NR prediction, as more PM corrections are included. Its accuracy is somewhat better than SEOBNRv5, despite the latter model using a Hamiltonian calibrated in the non-spinning ($a_{6}$) and spin-orbit coupling ($d_{\rm SO}$) sector (see below).